OpenStudy (anonymous):

Use the elimiation methjod to solve of quation 5u+3v=-18 2u+v=-5

OpenStudy (anonymous):

2u + v = -5 times 3 = 6u + 3v = -15 5u+3v=-18 - 6u+3u=-15 -------------- -1u = -3 u = 3 2u + v = -5 6 + v = -5 v = -11

OpenStudy (anonymous):

thanks